Sunday was a lovely sunny day after a cloudy start, ideal weather for going through the clubs hives to check all’s well.
the five main things we were checking , Queen cells, space, stores, disease, brood at all stages. Because there were so few of us we skipped some hives as they had either queen cells or virgin queens. Weather sunny, 18C with light wind.
Hive#1 Not inspected, awaiting laying new queen
Hive#2 Not inspected, awaiting laying of new queen
Hive#3 Not inspected, awaiting laying of new queen
Hive#4 Inspected by David P Brood Chamber (Double Deep) frames aligned the warm way. queen not seen(green 2014) Top Deep 20 queen cells , 9 frames brood, 2 stores, room ok. Bottom Deep 7 queen cells, 7 frames brood, 2 stores, room ok. Super 1 full, ready to take off, super 2 filling. Another super with foundation added under . Notes: A frame with one queen cell on and a frame of stores moved into a poly Nuclei. and another big queen cell used to replace a rubbish qc in Nuclei 15D( cardboard no 1) Temper: 10
Hive #5 Not gone through as sealed queen cell last week.
Hive #6 Inspected by Liz H Brood chamber (shallow over deep) Queen seen and marked blue., No QCs . 4 frames deep brood, none in shallow, 9 stores in super. Temper: 10
Hive# 7 Inspected by David P Two Shallow brood boxes. Queen not seen but marked previously.(blue) No QCs top brood box – 8 Frames, 7 frames in bottom box. 5 frames stores , 4 frames of room. one super, part filling. Temper: 7
Hive#8 Inspected by Liz H One Deep Brood box. and 3 supers Queen seen, No QCs, but brood all stages. 6 deep frames of brood, 3 stores, not much room in single deep. Notes: As little room , moved queen excluder up over bottom empty super . Temper: 7
Hive #9 Not gone through
15A wooden end nuc,qns. 2 brood half stores, b.a.s. lot of drone brood, slow making up.
15D cardboard no1, qns. 1 brood , 1 qc replaced from colony 4. stores ok.
15E cardboard no2 qns. old drone brood only, no eggs or larva. very sluggish bees.
15G cardboard no4 qns. 3 brood , 1 stores, strapped for space so transferred into Deep Brood box
15H Poly Nuc set up from hive #4 – no check required till 26/7